# UMoE: Unifying Attention and FFN with Shared Experts

### v1

@@ Abstract
Sparse Mixture of Experts (MoE) architectures have emerged as a promising approach for scaling Transformer models. While initial works primarily incorporated MoE into feed-forward network (FFN) layers, recent studies have explored extending the MoE paradigm to attention layers to enhance model performance. However, existing attention-based MoE layers require specialized implementations and demonstrate suboptimal performance compared to their FFN-based counterparts. In this paper, we aim to unify the MoE designs in attention and FFN layers by introducing a novel reformulation of the attention mechanism, revealing an underlying FFN-like structure within attention modules. Our proposed architecture, UMoE, achieves superior performance through attention-based MoE layers while enabling efficient parameter sharing between FFN and attention components.
@@ 1 Introduction
We identify two distinctions between attention-MoE and FFN-MoE implementations that likely account for the observed performance differential: (1) the different expert design between attention and FFN layers, and (2) attention-MoE’s necessity to compromise the expressiveness of vanilla attention mechanisms to accommodate sparse computation [12]. Motivated by these observations, we investigate a compelling question: can we reformulate attention to reveal an underlying structure compatible with the same expert design as FFN layers, without sacrificing the expressive power of the attention mechanism? This is a challenging question due to the inherent complexity of attention mechanisms, including multiple projections and softmax calculations, which fundamentally differ from the straightforward two-matrix multiplication pattern of FFNs.
@@ 2 Related Work
Several recent approaches have explored extending the MoE paradigm to attention layers in Transformers [12, 13], with a primary focus on expert design. Since attention layers lack the consecutive matrix multiplication pattern found in FFNs, these approaches necessitate expert designs that differ from FFN-MoE models. The Mixture-of-Attention (MoA) [12] propose to conceptualize individual attention heads as experts, scaling attention layers by increasing the number of attention heads. However, introducing sparsity into attention layers presents a significant challenge: query vectors computed by a specific expert (or head) require corresponding key and value vectors from the same expert, necessitating identical expert activation across all tokens. To address this constraint, MoA implements distinct query and output projections per head while maintaining shared key and value projections across attention heads.
@@
q = x @ W_q[i]
# K and V (the token embeddings X) are shared across experts.
@@
Figure 3: Implementation details of a UMoE layer. The input consists of a sequence X containing n token embeddings and x representing the final token embedding. For simplicity, this implementation focuses on computing the output for the last token.
@@ 3 Method
The attention mechanism is the core of Transformers [22], processing token embeddings to capture contextual relationships. However, its structure differs from FFN layers, which complicates the unification of MoE designs across both modules. In this section, we present two alternative formulations of attention, pre-mixing and post-mixing, that reveal an inherent FFN-like structure within attention layers. Based on these formulations, we introduce a novel MoE architecture, UMoE.
@@ 3.1 Formulations of Attention
Consider a sequence of token embeddings $\mathbf{X}\in\mathbb{R}^{n\times d}$, where $n$ is the sequence length and $d$ is the embedding dimension. In multi-head attention, each token attends to all other tokens in the sequence through query, key, and value projections. For a single token $\boldsymbol{x}$ (e.g., the last token in the sequence for simplicity), its attention output is computed as:
@@
  • Eq. 7: A new interpretation where token embeddings are first aggregated into contextualized representations, i.e., weighted averages of all tokens, before being processed by the value ($\mathbf{W}_{v}^{i}$) and output ($\mathbf{W}_{o}^{i}$) projections. We term this formulation as pre-mixing attention.
@@
In this formulation, token embeddings are transformed by two successive projections independently for each token, before being aggregated using the attention weights.
@@ 3.2 UMoE
Referring to Eq. 2, we observe that the primary distinction between FFN-MoE layers and UMoE-Att layers lies in their expert inputs: FFN experts operate on individual token embeddings $\boldsymbol{x}$, while attention experts process weighted combinations of all token embeddings. This reveals a relationship: FFN-MoE layers can be interpreted as a specialized case of pre-mixing attention layers where the attention matrix is constrained to an identity matrix, limiting each token to self-attention only.
@@
Both the experts in UMoE-Att and the FFN layers of Transformer consist of two consecutive matrices. While attention layer experts utilize a relatively small intermediate size ($d_{v}$), FFN layers typically employ larger dimensions. Recent advances in FFN-MoE models suggest the efficacy of using FFN layers with reduced intermediate sizes as experts [6, 15, 3]. This insight enables the direct adoption of experts in attention layers for FFN layers, resulting in a fully MoE architecture, denoted as UMoE. Fig. 3 presents the pseudo-code of a UMoE layer, where the MoE paradigm is applied to both FFN and attention layers using a shared expert set. Notably, to facilitate parameter sharing, experts are implemented as two-layer FFNs with an intermediate size of $d_{v}$ and incorporate a non-linear activation function between matrix multiplications.
@@
The token mixing operation in pre-mixing attention is a weighted summation over token embeddings, which can be implemented as vanilla attention, accepting $Q$, $K$, $V$ matrices as input and producing an output matrix. Each token generates distinct query vectors for different experts, while values (input token embeddings) and their associated keys are shared across experts. To generate expert-dependent queries for input tokens, each expert requires an additional query projection matrix. To mitigate the parameter count disparity with existing MoE models, where experts typically comprise two matrices, we employ low-rank matrices [23] for query projection within UMoE experts. For a given token $\boldsymbol{x}$, the query for expert $i$ is computed as:
@@
where the first term is shared across all experts, while the second term is expert-specific with unique parameters, $\mathbf{W}_{a}^{i}\in\mathbb{R}^{d\times r}$ and $\mathbf{W}_{b}^{i}\in\mathbb{R}^{r\times d_{k}}$, for each expert.
@@
As illustrated in Fig. 2, UMoE integrates three key components: (1) experts implemented as fine-grained FFNs with dual low-rank query projection matrices, (2) pre-mixing attention mechanism utilizing shared keys and values across experts, and (3) the top-k router for expert selection. It is noteworthy that while MoA [12] also shares keys and values across experts, the values of MoA are the results after applying a value linear transformation to the input token embeddings. In contrast, the values of UMoE directly refer to the input token embeddings.
@@ 3.3 Discussion
Vanilla attention and pre-mixing attention differ in terms of KV cache requirements and computational complexity. During inference, vanilla attention requires caching multiple keys and values per token, whereas pre-mixing attention requires only one key and token embedding per token. Regarding the computation, while vanilla attention performs weighted summation over low-dimensional value vectors, pre-mixing attention operates on input token embeddings, introducing a modest increase in computational complexity. This modest increase, however, becomes increasingly negligible as models scale to larger dimensions, effectively amortizing the additional computational overhead. A detailed comparative analysis is presented in Table LABEL:tab:complexity (A.1). Additionally, the abstract formulation of UMoE opens avenues for future research to explore more computationally efficient token mixing alternatives, such as linear attention mechanisms [24, 25].
@@ 4.1 Setup
Experimental Setup. UMoE is implemented as a decoder-only Transformer with rotary position embedding [34], following deepseek-MoE [6]. We evaluate two model configurations: The base models comprise $12$ layers with a hidden size of $768$, while the large models consist of $24$ layers with a hidden size of $2048$. These configurations yield dense models with 134134M and 1.11.1B parameters, respectively. MoE variants replace all attention or FFN layers with MoE layers. Due to computational constraints, unless otherwise specified, models are pretrained on 50B tokens from FineWeb-Edu with batch size of $1024$. For Wikitext-103, following Csordás et al. [13], models are trained for 100k steps, though models typically overfit within 20k steps. Detailed hyperparameters are provided in A.3.
@@ 4.2 Comparison with Baselines
Table 1: Comparison of Dense and Sparse Mixture-of-Experts (MoE) Models for Language Modeling. In MoE models, $A\times B$ denotes $B$ experts per layer with size AA in ‘#Total’ columns, while $B$ in ‘#Active’ columns indicates the number of experts activated per token. Gray entries in ‘#Total’ columns indicate shared parameters between attention and FFN modules in UMoE models. UMoE-Att refers to UMoE variants with MoE only applied to attention modules.
@@ 4.3 Ablations
Figure 5: Best valid PPL (top) and training loss (bottom) on Wikitext.
@@ 4.4 Expert Specialization
We also provide an analysis on the attention maps of UMoE in A.4, which confirms that higher-ranked experts show more focused attention distributions on relevant tokens compared to lower-ranked ones.
@@ 5 Conclusion
As for future work, we are looking at replacing the token mixing mechanism of UMoE with more efficient alternatives to enable scaling up the number of activated experts in attention layers. In addition, we are also interested in investigating architectures that unify attention and FFN into a single layer, given our finding that FFN layers function as a specialized case of attention with reduced expressiveness.
@@ A.1 Complexity
Figure 6: Implementation of a UMoE layer based on post-mixing attention. $X$ is a sequence of $n$ token embeddings. In the attention layer, tokens are processed by their top-$k$ experts independently. The output embeddings of all tokens are aggregated according to the attention weights.
@@ A.2 Post-mixing vs Pre-mixing
6 presents the pseudo-code of a UMoE layer based on post-mixing attention. We conducted preliminary experiments on Wikitext-103 and FineWeb-Edu datasets. As illustrated in 7, MoE models incorporating pre-mixing attention demonstrate substantially superior performance compared to their post-mixing counterparts.
As discussed in 3.1, these two reformulations of the attention mechanism are mathematically equivalent, owing to the absence of non-linear transformations within the matrix chain multiplication in attention heads. However, when grouping the value and output projections and implementing them as an FFN with a non-linear activation function, these formulations yield distinct outputs in the MoE layers. 8 illustrates MoE models implemented based on these two reformulations. Both approaches can be interpreted as extensions of conventional FFN-based MoE models. Specifically, the pre-mixing approach enhances FFN-based MoE models by contextualizing the inputs, while post-mixing attention enables MoE layers to incorporate other tokens’ outputs in generating the final output. Future research could explore the synergistic combination of post- and pre-mixing approaches to fully leverage contextual information.
@@
A.3 Hyperparameters
LABEL:tab:hyper_base and LABEL:tab:hyper give the parameters used for based and large models, respectively. MoA and SwitchHead utilize identical parameters as UMoE-Att, excluding the low-rank query projections. LABEL:tab:hyper_training details the hyperparameters used during training. For the Wikitext-103 dataset, we adopt the same hyperparameter configuration as Csordás et al. [13].
@@ A.4 Attention Analysis
9 presents attention maps for the final token prediction, revealing distinct patterns between higher and lower-ranked experts. Higher-ranked experts demonstrate more focused attention distributions that align with task requirements. For instance, in the Shakespeare question, Expert_0 and Expert_32 show pronounced attention weights on task-critical tokens "who wrote". Similarly, for the Tokyo question, the top expert exhibits sophisticated attention patterns by focusing on key contextual elements like "Japan" and "capital". These observations suggest that the routing mechanism effectively identifies experts capable of extracting task-relevant information through specialized attention patterns.
@@
To create a comprehensive visualization, we aggregated attention maps of all layers by summing them, as shown in 10. The values on the left (e.g., E0: 3.83) represent the sum of router scores received by experts at each rank position across all layers. The accumulated patterns reveal that higher-ranked experts (particularly E0 and E1) maintain more targeted attention distributions focused on question-relevant tokens, while lower-ranked experts tend to focus heavily on the initial token,

### v2

@@ Abstract
Sparse Mixture of Experts (MoE) architectures have emerged as a promising approach for scaling Transformer models. While initial works primarily incorporated MoE into feed-forward network (FFN) layers, recent studies have explored extending the MoE paradigm to attention layers to enhance model performance. However, existing attention-based MoE layers require specialized implementations and demonstrate suboptimal performance compared to their FFN-based counterparts. In this paper, we aim to unify MoE designs in attention and FFN layers by introducing a novel reformulation of the attention mechanism, that reveals an underlying FFN-like structure within attention modules. Our proposed architecture, UMoE, achieves superior performance through attention-based MoE layers while enabling efficient parameter sharing between FFN and attention components.
@@ 1 Introduction
We identify two distinctions between attention-MoE and FFN-MoE implementations that likely account for the observed performance differential: (1) the different expert design between attention and FFN layers, and (2) attention-MoE’s necessity to compromise the expressiveness of vanilla attention mechanisms to accommodate sparse computation [12]. Motivated by these observations, we investigate a compelling question: can we reformulate attention to reveal an underlying structure compatible with the same expert design as FFN layers, without compromising the expressive power of the attention mechanism? This is a challenging question due to the inherent complexity of attention mechanisms, including multiple projections and softmax calculations, which fundamentally differ from the straightforward two-matrix multiplication pattern of FFNs.
@@ 2 Related Work
Several recent approaches have explored extending the MoE paradigm to attention layers in Transformers [12, 13], with a primary focus on expert design. Because attention layers lack the consecutive matrix multiplication pattern found in FFNs, these approaches necessitate expert designs that differ from FFN-MoE models. The Mixture-of-Attention (MoA) [12] propose to conceptualize individual attention heads as experts, scaling attention layers by increasing the number of attention heads. However, introducing sparsity into attention layers presents a significant challenge: query vectors computed by a specific expert (or head) require corresponding key and value vectors from the same expert, necessitating identical expert activation across all tokens. To address this constraint, MoA implements distinct query and output projections per head while maintaining shared key and value projections across attention heads.
@@
q = q_shared + x @ W_a[i] @ W_b[i]
# K and V (the hidden states X) are shared across experts.
@@
Figure 3: Implementation details of a UMoE layer. The input consists of a sequence X containing n token hidden states and x representing the final hidden state. For simplicity, this implementation focuses on computing the output for the last token.
@@ 3 Method
The attention mechanism is the core of Transformers [22], processing token hidden states to capture contextual relationships. However, its structure differs from FFN layers, which complicates the unification of MoE designs across both modules. In this section, we present two alternative formulations of attention, pre-mixing and post-mixing, that reveal an inherent FFN-like structure within attention layers. Based on these formulations, we introduce a novel MoE architecture, UMoE.
@@ 3.1 Formulations of Attention
Consider a sequence of token hidden states $\mathbf{X}\in\mathbb{R}^{n\times d}$, where $n$ is the sequence length and $d$ is the hidden dimension. In multi-head attention, each token attends to all other tokens in the sequence through query, key, and value projections. For a single token $\boldsymbol{x}$ (e.g., the last token in the sequence for simplicity), its attention output is computed as:
@@
  • Eq. 7: A new interpretation where token hidden states are first aggregated into contextualized representations, i.e., weighted averages of all tokens, before being processed by the value ($\mathbf{W}_{v}^{i}$) and output ($\mathbf{W}_{o}^{i}$) projections. We term this formulation as pre-mixing attention.
@@
In this formulation, token hidden states are transformed by two successive projections independently for each token, before being aggregated using the attention weights.
@@ 3.2 UMoE
Referring to Eq. 2, we observe that the primary distinction between FFN-MoE layers and UMoE-Att layers lies in their expert inputs: FFN experts operate on individual token hidden states $\boldsymbol{x}$, while attention experts process weighted combinations of all token hidden states. This reveals a relationship: FFN-MoE layers can be interpreted as a specialized case of pre-mixing attention layers where the attention matrix is constrained to an identity matrix, limiting each token to self-attention only.
@@
Both the experts in UMoE-Att and the FFN layers of Transformer consist of two consecutive matrices. While attention layer experts utilize a relatively small intermediate size ($d_{v}$), FFN layers typically employ larger dimensions. Recent advances in FFN-MoE models suggest the efficacy of using FFN layers with reduced intermediate sizes as experts [6, 15, 3]. This insight enables the direct adoption of experts in attention layers for FFN layers, resulting in a fully MoE architecture, denoted as UMoE. Fig. 2 illustrates the architecture of a UMoE layer., where the MoE paradigm is applied to both FFN and attention layers using a shared expert set. Notably, to facilitate parameter sharing, experts are implemented as two-layer FFNs with an intermediate size of $d_{v}$ and incorporate a non-linear activation function between matrix multiplications.
@@
The token mixing operation in pre-mixing attention is a weighted summation over token hidden states, which can be implemented as vanilla attention, accepting $Q$, $K$, $V$ matrices as input and producing an output matrix. Each token generates distinct query vectors for different experts, while values (hidden states) and their associated keys are shared across experts. To generate expert-dependent queries for input tokens, each expert requires an additional query projection matrix, leading to a notable increase in parameters. To mitigate the parameter count disparity with existing MoE models, where experts typically comprise two matrices, we employ low-rank matrices [23] for query projection within UMoE experts. For a given token $\boldsymbol{x}$, the query for expert $i$ is computed as:
@@
where the first term is shared across all experts, while the second term is expert-specific with unique parameters, $\mathbf{W}_{a}^{i}\in\mathbb{R}^{d\times r}$ and $\mathbf{W}_{b}^{i}\in\mathbb{R}^{r\times d_{k}}$, for each expert. Fig. 3 presents the pseudo-code of a UMoE layer.
@@
As illustrated in Fig. 2, UMoE integrates three key components: (1) experts implemented as fine-grained FFNs with dual low-rank query projection matrices, (2) pre-mixing attention mechanism utilizing shared keys and values across experts, and (3) the top-k router for expert selection. It is noteworthy that while MoA [12] also shares keys and values across experts, the values of MoA are the results after applying a value linear transformation to the input token hidden states. In contrast, the values of UMoE directly refer to the input token hidden states.
@@ 3.3 Discussion
Vanilla attention and pre-mixing attention differ in terms of KV cache requirements and computational complexity. During inference, vanilla attention requires caching multiple keys and values per token, whereas pre-mixing attention requires only one key and hidden state per token. While grouped-query attention (GQA) is commonly adopted to reduce the KV cache size in attention layers, it cannot be combined with pre-mixing attention, since the latter already maintains only a single key–value pair per token. Instead, multi-head latent attention (MLA) [8], a promising alternative to GQA, can be applied by introducing a down-projection to the hidden states before the token-mixing operation. Regarding the computation, while vanilla attention performs weighted summation over low-dimensional value vectors, pre-mixing attention operates on input token hidden states, introducing a modest increase in computational complexity. This modest increase, however, becomes increasingly negligible as models scale to larger dimensions, effectively amortizing the additional computational overhead. A detailed comparative analysis is presented in Table LABEL:tab:complexity (A.1). Additionally, the abstract formulation of UMoE opens avenues for future research to explore more computationally efficient token mixing alternatives, such as linear attention mechanisms [24, 25].
@@ 4.1 Setup
Experimental Setup. UMoE is implemented as a decoder-only Transformer with rotary position embedding [34], following deepseek-MoE [6]. The load balancing loss proposed by Switch Transformer [5] is adopted to encourage a balanced load across experts. The experts in our experiments are implemented as two-layer MLPs to ensure fair comparison with the baselines. [^2] UMoE is designed as a flexible framework agnostic to the expert implementation. Using more advanced gated MLP variants, such as SwiGLU, may lead to improved performance, as the gating mechanism can enhance the expressiveness of both attention and feed-forward layers.
We evaluate two model configurations: The base models comprise $12$ layers with a hidden size of $768$, while the large models consist of $24$ layers with a hidden size of $2048$. These configurations yield dense models with 134134M and 1.11.1B parameters, respectively. MoE variants replace all attention or FFN layers with MoE layers. Due to computational constraints, unless otherwise specified, models are pretrained on 50B tokens from FineWeb-Edu with batch size of $1024$. For Wikitext-103, following Csordás et al. [13], models are trained for 100k steps, though models typically overfit within 20k steps. Detailed hyperparameters are provided in A.4.
@@ 4.2 Comparison with Baselines
Table 1: Comparison of Dense and Sparse Mixture-of-Experts (MoE) Models for Language Modeling. In MoE models, $A\times B$ denotes $B$ experts per layer with size $A$ in ‘#Total’ columns, while $B$ in ‘#Active’ columns indicates the number of experts activated per token. Gray entries in ‘#Total’ columns indicate shared parameters between attention and FFN modules in UMoE models. UMoE-Att refers to UMoE variants with MoE only applied to attention modules.
@@ 46 The, _The, _the, the, _Our _a, _his, _my, Your, _Your
Figure 5: Best valid PPL (top) and training loss (bottom) on Wikitext.
@@ 4.4 Expert Specialization
We also provide an analysis on the attention maps of UMoE in A.5, which confirms that higher-ranked experts show more focused attention distributions on relevant tokens compared to lower-ranked ones.
@@ 5 Conclusion
As for future work, we are looking at replacing the token mixing mechanism with more efficient alternatives to enable scaling up the number of activated experts in attention layers. In addition, we are also interested in investigating architectures that unify attention and FFN into a single layer, given our finding that FFN layers function as a specialized case of attention with reduced expressiveness. Exploring parameter sharing across different Transformer layers [35, 36] is also a promising direction.
@@ A.1 Complexity
Figure 6: Implementation of a UMoE layer based on post-mixing attention. $X$ is a sequence of $n$ token hidden states. In the attention layer, tokens are processed by their top-$k$ experts independently. The output embeddings of all tokens are aggregated according to the attention weights.
@@ A.2 Post-mixing vs Pre-mixing
Fig. 6 presents the pseudo-code of a UMoE layer based on post-mixing attention. We conducted preliminary experiments on Wikitext-103 and FineWeb-Edu datasets. As illustrated in Fig. 7, MoE models incorporating pre-mixing attention demonstrate substantially superior performance compared to their post-mixing counterparts.
As discussed in Section 3.1, these two reformulations of the attention mechanism are mathematically equivalent, owing to the absence of non-linear transformations within the matrix chain multiplication in attention heads. However, when grouping the value and output projections and implementing them as an FFN with a non-linear activation function, these formulations yield distinct outputs in the MoE layers. Fig. 8 illustrates MoE models implemented based on these two reformulations. Both approaches can be interpreted as extensions of conventional FFN-based MoE models. Specifically, the pre-mixing approach enhances FFN-based MoE models by contextualizing the inputs, while post-mixing attention enables MoE layers to incorporate other tokens’ outputs in generating the final output. Future research could explore the synergistic combination of post- and pre-mixing approaches to fully leverage contextual information.
@@
A.3 Wall-clock Efficiency Analysis
Table 9 presents a quantitative comparison of end-to-end training throughput and inference latency across various model architectures. All experiments were conducted on a single NVIDIA H100 GPU. Inference latency corresponds to the pre-filling time for sequences of 1024 tokens, while training throughput is measured in tokens per second.
Across both scales, all Mixture-of-Experts (MoE) variants exhibit comparable inference latency but remain slower than the dense baseline, despite their similar theoretical MAC counts. This discrepancy primarily arises from additional computational overheads introduced by expert routing and sparse expert execution. Among MoE models, SwitchHead and UMoE show slightly higher latency due to their architectural designs—SwitchHead applies two MoE layers per attention head, whereas UMoE employs two MoE layers within each Transformer block.
These results underscore a persistent challenge in MoE-based architectures: while their theoretical efficiency is well established, current hardware and software infrastructures (e.g., routing kernels, expert parallelization) are not yet fully optimized to realize these potential speedups. Consequently, our main paper reports MAC-based efficiency as the primary metric. We anticipate that advances in GPU kernel design and distributed parallelization will further mitigate the observed wall-clock inefficiencies in future implementations.
Table 9: Comparison of training throughput and inference latency across models. Inference latency denotes the pre-filling time (1024 tokens) measured on a single H100 GPU. All results are averaged over multiple runs.
Model Type Model Model Size Pre-filling Latency (s) Training Throughput (tokens/s)
Base Model GPT 134M 0.1376 126,508
FFN-MoE 535M 0.1974 74,415
MoA 525M 0.1843 90,394
SwitchHead 533M 0.2253 69,374
UMoE-Att 547M 0.2006 76,538
UMoE 540M 0.2197 71,461
Large Model GPT 1.1B 0.2950 11,634
FFN-MoE 3.8B 0.4080 9,799
MoA 3.6B 0.3624 12,321
SwitchHead 3.7B 0.4328 8,882
UMoE-Att 3.8B 0.4120 11,377
UMoE 3.6B 0.4392 9,194
A.4 Hyperparameters
Table LABEL:tab:hyper_base and LABEL:tab:hyper give the parameters used for based and large models, respectively. It takes roughly a week for pretraining base models on FineWeb-Edu datasets. MoA and SwitchHead utilize identical parameters as UMoE-Att, excluding the low-rank query projections. Table!LABEL:tab:hyper_training details the hyperparameters used during training. For the Wikitext-103 dataset, we adopt the same hyperparameter configuration as Csordás et al. [13].
@@ A.5 Attention Analysis
Fig. 9 presents attention maps for the final token prediction, revealing distinct patterns between higher and lower-ranked experts. Higher-ranked experts demonstrate more focused attention distributions that align with task requirements. For instance, in the Shakespeare question, Expert_0 and Expert_32 show pronounced attention weights on task-critical tokens "who wrote". Similarly, for the Tokyo question, the top expert exhibits sophisticated attention patterns by focusing on key contextual elements like "Japan" and "capital". These observations suggest that the routing mechanism effectively identifies experts capable of extracting task-relevant information through specialized attention patterns.
@@
To create a comprehensive visualization, we aggregated attention maps of all layers by summing them, as shown in Fig. 10. The values on the left (e.g., E0: 3.83) represent the sum of router scores received by experts at each rank position across all layers. The accumulated patterns reveal that higher-ranked experts (particularly E0 and E1) maintain more targeted attention distributions focused on question-relevant tokens, while lower-ranked experts tend to focus heavily on the initial token,